OpenStudy (anonymous):

Help me please!!! When looking at a rational function, Charles and Bobby have two different thoughts. Charles says that the function is defined at x = −2, x = 3, and x = 5. Bobby says that the function is undefined at those x values. Describe a situation where Charles is correct, and describe a situation where Bobby is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (anonymous):

When you plug in -2,3, or 5, you get 0/0, which is mathematical garbage, so we don't define the function at that point. However, carelessly dividing out the numerator and denominator without remembering the special cases of x=-2,3,5 will result in f(x)=1 everywhere including x=-2,3,5 which is not quite correct

OpenStudy (anonymous):

\[f(x)=\frac{(x+2)(x-3)(x-5)}{(x+2)(x-3)(x-5)}\]

OpenStudy (anonymous):

I am not understanding how to do it.

OpenStudy (anonymous):

when you plug in x= 16 for\[\frac{x-16}{x-16}\]what do you get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

You get\[\frac{0}{0}\]Which makes no sense at all. So we can't say that the function at x=16 is 1

OpenStudy (anonymous):

\[\frac{0}{0}\] is not equal to any number

OpenStudy (anonymous):

Yes I understand that

OpenStudy (anonymous):

Ok, so expanding on the idea that where a function has a place where there can be a \[\frac{0}{0}\]we can combine multiple of these things that produce \[\frac{0}{0}\]

OpenStudy (anonymous):

The answer will be..... When you plug in those values, it equals 0/0 which is undefined, which makes Bobby correct. To make Charles correct, you have to_____________. I dont know how Charles is.

OpenStudy (anonymous):

Charles could be correct by remembering that x=-2,3,5 the function evaluates to \[\frac{0}{0}\] and so knowing it's not right to say the function exists at those points, he should agree with Bobby that the function does not exist at those points

OpenStudy (anonymous):

Thank you!!!